OpenStudy (anonymous):

Solve the identity:

OpenStudy (anonymous):

(cos x + cos y)^2 + (sin x - sin y)^2 = 2 + 2cos(x + y)

OpenStudy (anonymous):

Not sure what the process for proving a trig equation is.

OpenStudy (anonymous):

cos^2x + cos^2y +2cosxcosy + sin^2x + sin^2y -2sinxsiny = cos^2x + sin^2x + cos^2y + sin^2y + 2cosxcosy - 2sinxsiny

OpenStudy (anonymous):

now [ cos^2x + sin^2x =1 ] hence we have = cos^2x + sin^2x + cos^2y + sin^2y + 2cosxcosy - 2sinxsiny = 1+1+ 2(cosxcosy - sinxsiny) =2+2cos(x+y) [ cos(x+y) = cosxcosy - sinxsiny ]